OpenStudy (anonymous):

Which of the following points is on the graph of the equation y = x^2 ? A. (-2, 4) B. (-5, 10) C. (5, 10) D. (4, 8)

OpenStudy (raden):

what do you think ?

OpenStudy (anonymous):

I have no clue

OpenStudy (raden):

just doing trial error one by one in options. let it is A (-2,4) here x = -2 and y = 4 now distribute that point into y = x^2, get 4 = (-2)^2 4 = 4 so, it is right that the point (-2,4) is on the graph of y = x^2

OpenStudy (raden):

got it ?